# Learning Disentangled Stain and Structural Representations for Semi-Supervised Histopathology Segmentation

### v1

@@ 1 Introduction
Semi-supervised learning (SSL) offers an effective strategy to improve segmentation performance by leveraging a small amount of labeled data alongside a larger pool of unlabeled samples. However, existing SSL methods often overlook the unique visual complexity of histopathological images, which arises from two entangled yet semantically distinct sources: staining variations and tissue morphology. In particular, Hematoxylin and Eosin (H&E) staining introduces substantial inter-sample color variation due to inconsistencies in staining protocols and scanning devices, while structural variations reflect biologically meaningful changes in tissue architecture, including malignant progression. Prior works in semi-supervised segmentation, such as pseudo-labeling Zhang et al. (2021), consistency regularization Sohn et al. (2020), and teacher-student frameworks Tarvainen and Valpola (2017); Yu et al. (2019); Pham et al. (2025), generally process the image holistically and fail to disentangle stain and structure cues. This lack of targeted representation learning can limit their effectiveness in histopathology, where isolating stain-specific and structure-specific information is crucial for robust and interpretable segmentation.
@@ 2.1 Preliminary
Problem Settings. Following standard semi-supervised segmentation settings, we denote the labeled set as $\mathcal{D}_{l}=\{(x_{i}^{l},y_{i}^{l})\}_{i=1}^{N}$, and a larger unlabeled set $\mathcal{D}_{u}=\{(x_{k}^{u})\}_{m=1}^{M}$ ($N$: labeled samples, $M$: unlabeled samples, $M\gg N$). Here, $x_{i}^{l}$ and $x_{k}^{u}$ are input images from the labeled and unlabeled sets, respectively, and $y_{i}^{l}$ is the corresponding one-hot groundtruth for labeled sample $x_{i}^{l}$.
@@ 3.1 Datasets
We evaluated our proposed method on two publicly available histopathological benchmarks. GlaS Sirinukunwattana et al. (2017) is a gland segmentation dataset for colorectal adenocarcinoma across various cancer stages, containing 165 images (37 benign, 48 malignant, and 80 test) with resolutions around $775\times 522$ or $589\times 453$. CRAG Graham et al. (2019) targets gland segmentation in colon histopathology, comprising 213 annotated images, mostly sized $1512\times 1516$. It is particularly challenging due to substantial variations in gland morphology and staining. Following prior works, 20% of the images were used as a fixed test set, and the remaining 80% were split into five folds for cross-validation on both datasets.
@@ 3.2 Implementation Details
All experiments were conducted on a single NVIDIA RTX 3060 GPU with Python 3.11, PyTorch 2.5, and CUDA 12.2. For both GlaS and CRAG datasets, we adopted DeepLabV3+ with a ResNet-101 backbone as the segmentation model. Models were trained for 80 epochs on GlaS and 120 epochs on CRAG until convergence, with all images resized to $256\times 256$.
To enhance generalization, we applied shared augmentations across branches, including random horizontal/vertical flips and rotations. In addition, each student received branch-specific augmentations. The Color Student was trained with color-focused transformations using a OneOf strategy: either random color jittering (with brightness, contrast, saturation set to 0.3 and hue to 0.05) or histogram matching Ferrero et al. (2024), which preserves structural integrity while increasing color robustness. In contrast, the Structure Student was augmented with elastic deformation Faryna et al. (2021), perturbing spatial structures while retaining color, thereby enhancing sensitivity to boundaries and textures.
We used AdamW with a learning rate of $1\times 10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=1\times 10^{-8}$, and a weight decay of 0.05. A batch size of 4 empirically yielded the best performance. During inference, the student model with the best validation performance is selected for generating predictions.
@@ 3.3 Comparison with State-of-the-arts
5% Supervised Baseline $72.31\pm 3.19$ $58.82\pm 3.43$ $63.87\pm 3.29$ $50.73\pm 2.78$
Mean Teacher (NeurIPS 2017) $77.23\pm 3.08$ $64.59\pm 3.34$ $71.71\pm 2.06$ $58.78\pm 2.35$
UAMT (MICCAI 2019) $79.43\pm 3.28$ $67.85\pm 3.75$ $72.01\pm 2.45$ $59.53\pm 2.30$
Fixmatch (NeurIPS 2020) $79.15\pm 1.75$ $67.66\pm 2.21$ $71.67\pm 3.36$ $58.83\pm 3.56$
CPS (CVPR 2021) $79.55\pm 2.01$ $67.84\pm 3.21$ $74.34\pm 2.19$ $61.36\pm 2.58$
CT (MIDL 2022) $79.86\pm 1.49$ $68.13\pm 3.64$ $71.84\pm 3.93$ $58.34\pm 4.79$
CCVC (CVPR 2023) $80.84\pm 1.75$ $68.88\pm 2.28$ $73.28\pm 1.87$ $60.48\pm 2.17$
CorrMatch (CVPR 2024) $79.85\pm 1.96$ $67.79\pm 2.79$ $69.08\pm 2.31$ $55.39\pm 2.88$
FDCL (AAAI-25 Bridge Program) $81.64\pm 1.08$ $70.15\pm 1.52$ $74.55\pm 1.51$ $61.93\pm 1.84$
@@
10% Supervised Baseline $75.64\pm 3.99$ $62.48\pm 4.11$ $71.80\pm 3.11$ $58.91\pm 3.20$
Mean Teacher (NeurIPS 2017) $79.68\pm 3.33$ $68.08\pm 4.11$ $75.36\pm 2.17$ $62.77\pm 2.61$
UAMT (MICCAI 2019) $84.05\pm 0.78$ $74.20\pm 1.65$ $75.48\pm 3.16$ $63.05\pm 3.42$
Fixmatch (NeurIPS 2020) $81.13\pm 2.05$ $69.40\pm 3.71$ $74.87\pm 3.20$ $62.30\pm 3.38$
CPS (CVPR 2021) $84.29\pm 0.44$ $73.89\pm 0.88$ $78.08\pm 1.32$ $66.11\pm 1.79$
CT (MIDL 2022) $82.67\pm 1.15$ $71.65\pm 2.98$ $73.68\pm 1.86$ $60.44\pm 2.19$
CCVC (CVPR 2023) $83.78\pm 2.31$ $73.52\pm 2.99$ $74.97\pm 1.40$ $62.30\pm 1.76$
CorrMatch (CVPR 2024) $83.27\pm 2.23$ $72.59\pm 2.01$ $74.90\pm 1.41$ $61.93\pm 2.35$
FDCL (AAAI-25 Bridge Program) $84.35\pm 1.12$ $74.45\pm 1.14$ $76.28\pm 2.29$ $63.92\pm 2.88$
@@
Table 1: Quantitative results on GlaS-2017 and CRAG-2019 datasets on two labeled ratios. We highlight the best performance for each metric in bold.
@@ 3.4 Quantitative Results
As shown in Table 1, CSDS consistently outperforms previous SOTA methods under limited supervision on both GlaS and CRAG datasets. On the Glas dataset, with only 10% labeled data, CSDS achieves a Dice score of $82.86\pm 1.24$ and a Jaccard index of $72.01\pm 1.81$, which surpasses strong co-training baselines like FDCL and CPS. In 5% labeled ratio, CSDS narrows the Dice gap to full supervision to just 0.81%, while significantly outperforming teacher-student methods (e.g., Mean Teacher) and self-training methods (e.g., FixMatch). CSDS generalizes well across datasets, as evidenced by consistently outperforming all compared methods under all settings on CRAG, which presents more irregular and challenging nuclei structures.
@@ 3.6 Ablation study
0.05 Alternate $81.80\pm 2.38$ $70.44\pm 3.19$
Best student only $81.60\pm 1.88$ $70.13\pm 2.56$
@@
0.10 Alternate $84.56\pm 1.40$ $74.30\pm 1.93$
Best student only $84.34\pm 1.64$ $74.03\pm 2.28$
@@ 4 Conclusion
In this work, we introduced Color-Structure Dual-Student (CSDS), a novel semi-supervised segmentation framework tailored for histopathology images. By explicitly decoupling color and structural information into two specialized student networks, guided by a shared teacher model, CSDS effectively captures domain-specific variations. Our uncertainty-aware training strategy further improves pseudo-label reliability through color-adaptive and structure-adaptive weighting. Extensive experiments on GlaS and CRAG datasets under limited-label regimes demonstrate that CSDS outperforms strong baselines and recent state-of-the-art methods. These findings underscore the effectiveness of explicitly modeling visual cues unique to histopathology in semi-supervised segmentation.

### v2

@@ 1 Introduction
Semi-supervised learning (SSL) offers an effective strategy to improve segmentation performance by leveraging a small amount of labeled data alongside a larger pool of unlabeled samples. However, existing SSL methods often overlook the unique visual complexity of histopathological images, which arises from two entangled yet semantically distinct sources: staining variations and tissue morphology. In particular, Hematoxylin and Eosin (H&E) staining introduces substantial inter-sample color variation due to inconsistencies in staining protocols and scanning devices, while structural variations reflect biologically meaningful changes in tissue architecture, including malignant progression. Prior works in semi-supervised segmentation, such as pseudo-labeling, consistency regularization Sohn et al. (2020); Le et al. (2025), and teacher-student frameworks Tarvainen and Valpola (2017); Yu et al. (2019); Pham et al. (2025), generally process the image holistically and fail to disentangle stain and structure cues. This lack of targeted representation learning can limit their effectiveness in histopathology, where isolating stain-specific and structure-specific information is crucial for robust and interpretable segmentation.
@@ 2.1 Preliminary
Problem Settings. Following standard semi-supervised segmentation settings, we denote the labeled set as $\mathcal{D}_{l}=\{(x_{i}^{l},y_{i}^{l})\}_{i=1}^{N}$, and a larger unlabeled set as $\mathcal{D}_{u}=\{(x_{k}^{u})\}_{k=1}^{M}$ ($N$: labeled samples, $M$: unlabeled samples, $M\gg N$). Here, $x_{i}^{l}$ and $x_{k}^{u}$ are input images from the labeled and unlabeled sets, respectively, and $y_{i}^{l}$ is the corresponding one-hot groundtruth for labeled sample $x_{i}^{l}$. This setup formalizes the standard semi-supervised segmentation problem under a limited-label regime.
@@ 3.1 Datasets
We evaluated our proposed method on two publicly available histopathological benchmarks. GlaS Sirinukunwattana et al. (2017) is a gland segmentation dataset for colorectal adenocarcinoma across various cancer stages, containing 165 images (37 benign, 48 malignant, and 80 test) with resolutions around 775×522775$\times$ 522775 × 522 or 589×453589$\times$ 453589 × 453. CRAG Graham et al. (2019) targets gland segmentation in colon histopathology, comprising 213 annotated images, mostly sized 1512×15161512$\times$ 15161512 × 1516. It is particularly challenging due to substantial variations in gland morphology and staining. Following prior works, 20% of the images were used as a fixed test set, and the remaining 80% were split into five folds for cross-validation on both datasets.
@@ 3.2 Implementation Details
All experiments were conducted on a single NVIDIA RTX 3060 GPU (16 GB VRAM) using Python 3.11, PyTorch 2.5, and CUDA 12.2. We adopted DeepLabV3+ with a ResNet-101 backbone as the segmentation model for both GlaS and CRAG datasets, training for 80 and 120 epochs respectively, with all images resized to $\times 256$ 256.
To enhance generalization, we applied shared augmentations across branches, including random flips and rotations. Each student also received branch-specific augmentations tailored to its objective. The Color Student used either color jittering or histogram matching Ferrero et al. (2024) to improve robustness to appearance variations while preserving structure. In contrast, the Structure Student was augmented with elastic deformation Faryna et al. (2021) to better capture structural distortions and fine-grained morphology.
Models were optimized using AdamW ($\text{lr}=1\times 10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=1\times 10^{-8}$, weight decay = 0.05) with a batch size of 4. During inference, the student model with the best validation performance was used for prediction.
@@ 3.3 Comparison with State-of-the-arts
5% Supervised Baseline 72.31±3.1972.31$\pm$ 3.1972.31 ± 3.19 58.82±3.4358.82$\pm$ 3.4358.82 ± 3.43 63.87±3.2963.87$\pm$ 3.2963.87 ± 3.29 50.73±2.7850.73$\pm$ 2.7850.73 ± 2.78
Mean Teacher (NeurIPS 2017) 77.23±3.0877.23$\pm$ 3.0877.23 ± 3.08 64.59±3.3464.59$\pm$ 3.3464.59 ± 3.34 71.71±2.0671.71$\pm$ 2.0671.71 ± 2.06 58.78±2.3558.78$\pm$ 2.3558.78 ± 2.35
UAMT (MICCAI 2019) 79.43±3.2879.43$\pm$ 3.2879.43 ± 3.28 67.85±3.7567.85$\pm$ 3.7567.85 ± 3.75 72.01±2.4572.01$\pm$ 2.4572.01 ± 2.45 59.53±2.3059.53$\pm$ 2.3059.53 ± 2.30
Fixmatch (NeurIPS 2020) 79.15±1.7579.15$\pm$ 1.7579.15 ± 1.75 67.66±2.2167.66$\pm$ 2.2167.66 ± 2.21 71.67±3.3671.67$\pm$ 3.3671.67 ± 3.36 58.83±3.5658.83$\pm$ 3.5658.83 ± 3.56
CPS (CVPR 2021) 79.55±2.0179.55$\pm$ 2.0179.55 ± 2.01 67.84±3.2167.84$\pm$ 3.2167.84 ± 3.21 74.34±2.1974.34$\pm$ 2.1974.34 ± 2.19 61.36±2.5861.36$\pm$ 2.5861.36 ± 2.58
CT (MIDL 2022) 79.86±1.4979.86$\pm$ 1.4979.86 ± 1.49 68.13±3.6468.13$\pm$ 3.6468.13 ± 3.64 71.84±3.9371.84$\pm$ 3.9371.84 ± 3.93 58.34±4.7958.34$\pm$ 4.7958.34 ± 4.79
CCVC (CVPR 2023) 80.84±1.7580.84$\pm$ 1.7580.84 ± 1.75 68.88±2.2868.88$\pm$ 2.2868.88 ± 2.28 73.28±1.8773.28$\pm$ 1.8773.28 ± 1.87 60.48±2.1760.48$\pm$ 2.1760.48 ± 2.17
CorrMatch (CVPR 2024) 79.85±1.9679.85$\pm$ 1.9679.85 ± 1.96 67.79±2.7967.79$\pm$ 2.7967.79 ± 2.79 69.08±2.3169.08$\pm$ 2.3169.08 ± 2.31 55.39±2.8855.39$\pm$ 2.8855.39 ± 2.88
FDCL (AAAI-25 Bridge Program) 81.64±1.0881.64$\pm$ 1.0881.64 ± 1.08 70.15±1.5270.15$\pm$ 1.5270.15 ± 1.52 74.55±1.5174.55$\pm$ 1.5174.55 ± 1.51 61.93±1.8461.93$\pm$ 1.8461.93 ± 1.84
@@
10% Supervised Baseline 75.64±3.9975.64$\pm$ 3.9975.64 ± 3.99 62.48±4.1162.48$\pm$ 4.1162.48 ± 4.11 71.80±3.1171.80$\pm$ 3.1171.80 ± 3.11 58.91±3.2058.91$\pm$ 3.2058.91 ± 3.20
Mean Teacher (NeurIPS 2017) 79.68±3.3379.68$\pm$ 3.3379.68 ± 3.33 68.08±4.1168.08$\pm$ 4.1168.08 ± 4.11 75.36±2.1775.36$\pm$ 2.1775.36 ± 2.17 62.77±2.6162.77$\pm$ 2.6162.77 ± 2.61
UAMT (MICCAI 2019) 84.05±0.7884.05$\pm$ 0.7884.05 ± 0.78 74.20±1.6574.20$\pm$ 1.6574.20 ± 1.65 75.48±3.1675.48$\pm$ 3.1675.48 ± 3.16 63.05±3.4263.05$\pm$ 3.4263.05 ± 3.42
Fixmatch (NeurIPS 2020) 81.13±2.0581.13$\pm$ 2.0581.13 ± 2.05 69.40±3.7169.40$\pm$ 3.7169.40 ± 3.71 74.87±3.2074.87$\pm$ 3.2074.87 ± 3.20 62.30±3.3862.30$\pm$ 3.3862.30 ± 3.38
CPS (CVPR 2021) 84.29±0.4484.29$\pm$ 0.4484.29 ± 0.44 73.89±0.8873.89$\pm$ 0.8873.89 ± 0.88 78.08±1.3278.08$\pm$ 1.3278.08 ± 1.32 66.11±1.7966.11$\pm$ 1.7966.11 ± 1.79
CT (MIDL 2022) 82.67±1.1582.67$\pm$ 1.1582.67 ± 1.15 71.65±2.9871.65$\pm$ 2.9871.65 ± 2.98 73.68±1.8673.68$\pm$ 1.8673.68 ± 1.86 60.44±2.1960.44$\pm$ 2.1960.44 ± 2.19
CCVC (CVPR 2023) 83.78±2.3183.78$\pm$ 2.3183.78 ± 2.31 73.52±2.9973.52$\pm$ 2.9973.52 ± 2.99 74.97±1.4074.97$\pm$ 1.4074.97 ± 1.40 62.30±1.7662.30$\pm$ 1.7662.30 ± 1.76
CorrMatch (CVPR 2024) 83.27±2.2383.27$\pm$ 2.2383.27 ± 2.23 72.59±2.0172.59$\pm$ 2.0172.59 ± 2.01 74.90±1.4174.90$\pm$ 1.4174.90 ± 1.41 61.93±2.3561.93$\pm$ 2.3561.93 ± 2.35
FDCL (AAAI-25 Bridge Program) 84.35±1.1284.35$\pm$ 1.1284.35 ± 1.12 74.45±1.1474.45$\pm$ 1.1474.45 ± 1.14 76.28±2.2976.28$\pm$ 2.2976.28 ± 2.29 63.92±2.8863.92$\pm$ 2.8863.92 ± 2.88
@@
Table 1: Quantitative results on GlaS-2017 and CRAG-2019 datasets under two labeled ratios. We highlight the best performance for each metric in bold, and the second-best in underline.
@@ 3.4 Quantitative Results
As shown in Table 1, CSDS consistently outperforms previous SOTA methods under limited supervision on both GlaS and CRAG datasets. On the Glas dataset, with only 10% labeled data, CSDS achieves a Dice score of 82.86±1.2482.86$\pm$ 1.2482.86 ± 1.24 and a Jaccard index of 72.01±1.8172.01$\pm$ 1.8172.01 ± 1.81, which surpasses strong co-training baselines like FDCL and CPS. In 5% labeled ratio, CSDS narrows the Dice gap to full supervision to just 0.81%, while significantly outperforming teacher-student methods (e.g., Mean Teacher) and self-training methods (e.g., FixMatch). CSDS generalizes well across datasets, as evidenced by consistently outperforming all compared methods under all settings on CRAG, which presents more irregular and challenging nuclei structures.
@@ 3.6 Ablation study
0.05 Alternate 81.80±2.3881.80$\pm$ 2.3881.80 ± 2.38 70.44±3.1970.44$\pm$ 3.1970.44 ± 3.19
Best student only 81.60±1.8881.60$\pm$ 1.8881.60 ± 1.88 70.13±2.5670.13$\pm$ 2.5670.13 ± 2.56
@@
0.10 Alternate 84.56±1.4084.56$\pm$ 1.4084.56 ± 1.40 74.30±1.9374.30$\pm$ 1.9374.30 ± 1.93
Best student only 84.34±1.6484.34$\pm$ 1.6484.34 ± 1.64 74.03±2.2874.03$\pm$ 2.2874.03 ± 2.28
@@ 4 Conclusion
In this work, we introduced Color-Structure Dual-Student (CSDS), a novel semi-supervised segmentation framework tailored for histopathology images. By decoupling color and structural cues into two specialized student networks guided by a shared teacher, CSDS captures domain-specific variations more effectively. An uncertainty-aware training strategy further improves pseudo-label reliability via color- and structure-adaptive weighting. Experiments on GlaS and CRAG under limited-label settings show that CSDS consistently outperforms strong baselines and recent state-of-the-art methods, highlighting the benefit of explicitly modeling histopathology-specific visual cues.
Acknowledgments and Disclosure of Funding
This research was funded by NAFOSTED, Vietnam National Foundation for Science and Technology Development, grant number IZVSZ2_229539.